Write the point slope form of the equation of the line that passes through the point (-1,5) and has a slope of -1
Answer by rothauserc(4718)

The point slope form of a line is
:
(y - y1) = m(x - x1), where m is the slope and (x1, y1) is the point
:
(y - 5) = -(x + 1)
: